# The broader spectrum of in-context learning

### v2

@@ Abstract
In recent years, there has been substantial excitement about meta-learning (Finn et al., 2017; Wang, 2021). A notable substream of this work focuses on memory-based meta-learning (Santoro et al., 2016; Wang et al., 2016; Ortega et al., 2019), in which the inner learning process occurs solely within the memory (or activations) of a network. This excitement has been reignited by the recent demonstrations of In-Context Learning (ICL) of tasks from a few examples in transformer language models (Brown et al., 2020). These in-context learning abilities have been suggested to emerge from particular properties of the training data, such as bursty patterns in the training sequences (Chan et al., 2022a) or repeated parallel structures (Chen et al., 2024).
There has been substantial interest in studying ICL from theoretical and mechanistic perspectives (Chan et al., 2022a; Xie et al., 2022; Von Oswald et al., 2023; Pan et al., 2023; Akyürek et al., 2022; Singh et al., 2024; Zhang et al., 2024; Swaminathan et al., 2024; Frei and Vardi, 2024; Raventós et al., 2024; Park et al., 2024). However, these works have generally focused on few-shot supervised ICL—where the goal is to learn an input-output function from a few supervised input-output examples.
Yet language models perform many other kinds of in-context adaptation—such as adapting to tasks from instructions or role-play (Reynolds and McDonell, 2021; Shanahan et al., 2023), extrapolating time series (Gruver et al., 2024), or more active exploration and learning in context (e.g. Coda-Forno et al., 2023; Lampinen et al., 2024)—that do not cleanly fit within the few-shot supervised ICL paradigm. Likewise, there is a much broader meta-learning literature studying much more complex types of meta-learned in-context learning (e.g. Wang et al., 2021; Laskin et al., 2023; Bauer et al., 2023).
Thus, there is a gap between our theoretical and mechanistic understanding of ICL in language models—which tends to focus on few-shot supervised ICL—and the more general capabilities for in-context adaptation that they exhibit in practice. Are these qualitatively different phenomena, or can we understand them as part of an overarching meta-learning process? This note is intended to bridge this gap, and discusses some of the consequences.
See Fig. 1 for an overview of our argument. Our goal is to communicate the following perspective:
@@ 1 Background
Likewise, language models are not merely few-shot supervised learners—they are capable of much more general in-context adaptation from other kinds of in-context cues—such as instructions (e.g. Fig. 1 in Ouyang et al., 2022), explanations (Lampinen et al., 2022), or role play (Reynolds and McDonell, 2021; Shanahan et al., 2023). They also exhibit more general sequence capabilities (Gruver et al., 2024) and even hierarchical adaptation (Coda-Forno et al., 2023). Thus, the parallels between the types of adaptation considered in the meta-learning and goal-conditioning literature, and the in-context adaptation capabilities of language models, go far beyond the few-shot supervised ICL regime.
@@ 3 Standard Few-Shot Supervised ICL
Seen from this perspective, supervised ICL resembles certain kinds of analogical reasoning (cf. Gentner, 1983; Holyoak, 2012). For example, a classic analogy problem like Paris : France :: London : ? requires recognizing the relation on the left (is the capital of), and then applying a higher-order relation (that the relations on each side should be the same) to infer the missing element on the right.
@@ 5 Roots of ICL in language processing
In the previous section, we highlighted some of the sophisticated forms of in-context learning that language models exhibit. Where do these capabilities originate? Chan et al. (2022a) highlighted how standard few-shot classification ICL could emerge from simple statistical properties—burstiness and long tailed distributions—that are known properties of language. Likewise, Chen et al. (2024) study how parallel structures support learning of ICL. Here, we correspondingly suggest that the broader spectrum of in-context learning capabilities arises analogously from the multi-scale sequential dependency structures in language datasets, and the need to adapt to the long tail of information in the training corpus. In this section, we illustrate this perspective with examples connecting various kinds of ICL to linguistic properties.
@@ 6 Generalization of ICL
Figure 2: Examples of learning a function from examples in context, and then using that knowledge more flexibly than the initial examples would suggest. (This chat is edited from a real interaction with an LM; the full transcript is in Appx. B.)
@@ 6.1 Learning something novel
Many prior works studying the theoretical basis of in-context learning have already evaluated some kind of generalization to learning novel tasks in context, within the scope of their paradigms. For example, Xie et al. (2022) considered generalization under distribution shift, and Chan et al. (2022a) evaluated few-shot learning of novel categories. Likewise, theoretical studies of in-context learning of linear regression tasks (e.g. Zhang et al., 2024) have considered how ICL generalizes to learning when the function to regress (or the shot inputs) are sampled from different distributions. However, as our theoretical understanding of in-context learning expands, we hope that studies will consider generalization beyond the FSS-ICL paradigm. Some works have already explored this in some cases, for example, Ramesh et al. (2024) study compositional generalization to function compositions specified through explicit instructions rather than through examples. However, it is also interesting to consider generalization across different methods by which information can be presented.
@@ 6.2 Learning in varied formats
As we have outlined above, language models can learn the same task in context from many different types of cues. It is therefore interesting to understand the extent to which models perform similarly or differently across different ways of presenting the task. There has been some investigation of models variability in FSS-ICL performance from minor changes to the format or ordering of the prompt examples (e.g. Lu et al., 2022; Sclar et al., 2023). However, we believe there should be more systematic investigation of other types of ICL.
@@ 7 Discussion
Transfer and/or shared circuitry between different kinds of ICL: One intriguing possibility is that there may be transfer or shared circuitry between the way models implement different kinds of ICL—analogously to how language models may learn more complex syntactic structures by piecing together what they learn from related simpler ones (Misra and Mahowald, 2024). For example, whether a function is learned from examples or instructions, the application behavior is the same—and thus the circuitry for applying what is learned in context might overlap for the two cases. Similarly, copying circuitry used for more sophisticated ICL (e.g. in induction heads, Olsson et al., 2022) might be learned in part from predicting simpler single-token repetitions, or parallel syntactic structures. In turn, these linguistic structures, and the circuits that support them, might also contribute to more complex kinds of ICL. For example, the use of parallel relations in syntactic structures or FSS-ICL might relate to model performance on analogy tasks (Webb et al., 2023); [^6] Though cf. Lewis and Mitchell (2024) and Webb et al. (2024) for some further debate. or more complex types of ICL like meta-learning in context (Coda-Forno et al., 2023). Considering these interactions would therefore be important to understanding model behaviors and internal mechanisms more completely.
Interference between different kinds of ICL: Taking a broader perspective on ICL may provide a useful perspective on some instances where models fail to learn the intended task in-context, e.g. ignoring flipped labels (Min et al., 2022; Wei et al., 2023), or fail to correctly integrate across interactions in context (Qiu et al., 2024); these may be due to interference from other kinds of in-context or in-weights learning mechanisms (e.g. more basic associations of labels to particular tasks). Likewise, effects of seemingly-irrelevant factors like prompt formatting or ordering (e.g. Sclar et al., 2023; Lu et al., 2022) may stem in part from models learning about how these features predict task contexts across the training distribution. There may also be dynamic effects over training, in which different types of ICL dominate at different times (Singh et al., 2024; Park et al., 2024).

### v3

@@ Abstract
In recent years, there has been substantial excitement about meta-learning (Finn et al., 2017; Wang, 2021). A notable substream of this work focuses on memory-based meta-learning (Santoro et al., 2016; Wang et al., 2016; Ortega et al., 2019), in which the inner learning process occurs solely within the memory (or activations) of a network. This excitement has been reignited by the recent demonstrations of In-Context Learning (ICL) of tasks from a few examples in transformer language models (Brown et al., 2020; Dong et al., 2022). These in-context learning abilities have been suggested to emerge from particular properties of the training data, such as bursty patterns in the training sequences (Chan et al., 2022a) or repeated parallel structures (Chen et al., 2024).
There has been substantial interest in studying ICL from theoretical and mechanistic perspectives (Xie et al., 2022; Von Oswald et al., 2023; Pan et al., 2023; Akyürek et al., 2022; Wang et al., 2023; Zhang et al., 2024; Swaminathan et al., 2024; Raventós et al., 2024). However, the majority of these works have focused on few-shot supervised ICL—where the goal is to learn an input-output function from a few supervised input-output examples.
Yet language models perform many other kinds of in-context adaptation—such as adapting to tasks from instructions or role-play (Reynolds and McDonell, 2021; Shanahan et al., 2023), extrapolating time series (Gruver et al., 2024), or more active exploration and learning in context (e.g. Coda-Forno et al., 2023; Lampinen et al., 2024)—that do not cleanly fit within the few-shot supervised ICL paradigm. Likewise, there is a much broader literature studying more complex types of meta-learned in-context learning (e.g. Wang et al., 2021; Laskin et al., 2023; Bauer et al., 2023).
Thus, there is a gap between our theoretical and mechanistic understanding of ICL in language models—which tends to focus on few-shot supervised ICL—and the more general capabilities for in-context adaptation that they exhibit in practice. Are these qualitatively different phenomena, or can we understand them as part of an overarching meta-learning process? This paper is intended to bridge this gap, and discuss some of the consequences. See Fig. 1 for an overview of our argument.
Our goal is to communicate the following perspective:
@@ 1 Background
Likewise, language models are not merely few-shot supervised learners—they are capable of much more general in-context adaptation from other kinds of in-context cues—such as instructions (e.g. Fig. 1 in Ouyang et al., 2022), explanations (Lampinen et al., 2022), role play (Reynolds and McDonell, 2021; Shanahan et al., 2023), and other kinds of prompting (Schulhoff et al., 2024). They also exhibit more general sequence capabilities (Gruver et al., 2024) and even hierarchical adaptation (Coda-Forno et al., 2023). Thus, the parallels between the types of adaptation considered in the meta-learning and goal-conditioning literature, and the in-context adaptation capabilities of language models, go far beyond the few-shot supervised ICL regime.
@@ 3 Standard Few-Shot Supervised ICL
Seen from this perspective, supervised ICL resembles certain kinds of analogical reasoning (cf. Gentner, 1983; Holyoak, 2012)—as prior works have noted (Dong et al., 2022). For example, a classic analogy problem like Paris : France :: London : ? requires recognizing the relation on the left (is the capital of), and then applying a higher-order relation (that the relations on each side should be the same) to infer the missing element on the right.
@@ 5 Roots of ICL in language processing
In the previous section, we highlighted some of the sophisticated forms of in-context learning that language models exhibit. Where do these capabilities originate? Chan et al. (2022a) highlighted how standard few-shot classification ICL could emerge from simple statistical properties—burstiness and long tailed distributions—that are known properties of language. Likewise, Chen et al. (2024) study how parallel structures support learning of ICL. Other works have highlighted how more complex contextual structures in pretraining may contribute to in-context learning (Han et al., 2023). Here, we correspondingly suggest that the broader spectrum of in-context learning capabilities arises analogously from the multi-scale sequential dependency structures in language datasets, and the need to adapt to the long tail of information in the training corpus. In this section, we illustrate this perspective with examples connecting various kinds of ICL to linguistic properties.
@@ 6.1 Learning something novel
Many prior works studying the theoretical basis of in-context learning have already evaluated some kind of generalization to learning novel tasks in context, within the scope of their paradigms. For example, Xie et al. (2022) considered generalization under distribution shift, and Chan et al. (2022a) evaluated few-shot learning of novel categories. Likewise, theoretical studies of in-context learning of linear & nonlinear regression tasks (e.g. Zhang et al., 2024; Raventós et al., 2024; Frei and Vardi, 2024) have considered how ICL generalizes to learning when the function to regress (or the shot inputs) are sampled from different distributions. However, as our theoretical understanding of in-context learning expands, we hope that studies will consider generalization beyond the FSS-ICL paradigm. Some works have already explored this in some cases, for example, Ramesh et al. (2024) study compositional generalization to function compositions specified through explicit instructions rather than through examples. However, it is also interesting to consider generalization across different methods by which information can be presented.
@@ 6.2 Learning in varied formats
As we have outlined above, language models can learn the same task in context from many different types of cues. It is therefore interesting to understand the extent to which models perform similarly or differently across different ways of presenting the task. There has been some investigation of models variability in FSS-ICL performance from minor changes to the format or ordering of the prompt examples (e.g. Lu et al., 2022; Sclar et al., 2023). However, we believe there should be more systematic investigation of other types of ICL. We show some preliminary examples of such analyses in Appendix LABEL:appx:explanations_experiments—showing similarities and differences in how a language model adapts to tasks from prompts containing instructions, explanations and examples.
@@
Figure 2: Examples of learning a function from examples in context, and then using that knowledge more flexibly than the initial examples would suggest. (Edited from a real interaction with an LM; the full transcript is in Appx. B.)
@@ 7 Discussion
Transfer and/or shared circuitry between different kinds of ICL: One intriguing possibility is that there may be transfer or shared circuitry between the way models implement different kinds of ICL—analogously to how language models may learn more complex syntactic structures by piecing together what they learn from related simpler ones (Misra and Mahowald, 2024). For example, whether a function is learned from examples or instructions, the application behavior is the same—and thus the circuitry for applying what is learned in context might overlap for the two cases. Indeed, one recent work finds that although the function-vector heads identified using instructions or examples are mostly distinct, there is a subset that overlap (Davidson et al., 2025). Similarly, copying circuitry used for more sophisticated ICL (e.g. in induction heads, Olsson et al., 2022) might be learned in part from predicting simpler single-token repetitions, or parallel syntactic structures. In turn, these linguistic structures, and the circuits that support them, might also contribute to more complex kinds of ICL. For example, the use of parallel relations in syntactic structures or FSS-ICL might relate to model performance on analogy tasks (Webb et al., 2023); [^6] Though cf. Lewis and Mitchell (2024) and Webb et al. (2024) for some further debate. or more complex types of ICL like meta-learning in context (Coda-Forno et al., 2023). Conversely, if models can use multiple distinct strategies for an ICL behavior (cf. Bhattamishra et al., 2024), then the underlying mechanisms might interact in complex ways. Indeed, several recent works show complex patterns of cooperation and competition between different types of ICL in overlapping circuits (Singh et al., 2025), and transitions between different mechanistic implementations of ICL over the course of training (Park et al., 2024; Yin and Steinhardt, 2025). Considering these interactions would therefore be important to understanding model behaviors and internal mechanisms more completely.
Interference between different kinds of ICL: Taking a broader perspective on ICL may provide a useful perspective on some instances where models fail to learn the intended task in-context. For example, this includes e.g. ignoring flipped labels (Min et al., 2022; Wei et al., 2023) or instructions (Webson and Pavlick, 2022), or overly-fixating on label patterns (Yan et al., 2024), or failing to correctly integrate across interactions in context (Qiu et al., 2024). These failures may be due to interference from other kinds of in-context or in-weights learning mechanisms (e.g. more basic associations of labels to particular tasks). Likewise, effects of seemingly-irrelevant factors like prompt formatting or ordering (e.g. Sclar et al., 2023; Lu et al., 2022) may stem in part from models learning about how these features predict task contexts across the training distribution. There may also be dynamic effects over training, in which different types of ICL dominate at different times (Singh et al., 2024; Park et al., 2024).